OpenStudy (leyla.tas123):

what is the length of each side of a square if the A=49 m^2

OpenStudy (chpatterson):

www.algebra.com › Surface area

OpenStudy (rcp031):

A square has equal lengths L*L=L^2=7^2=49.

OpenStudy (leyla.tas123):

thank you so much both of you!

OpenStudy (chpatterson):

np